# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 4.2. Parabolic restriction
In particular, the tensor product $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}:=\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}\mathrm{U}_{\hbar}(\mathfrak{g})$ has a well-defined algebra structure. Denote by $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})^{\mathrm{gen}}$ the corresponding generic version of Harish-Chandra bimodules for $\mathrm{GL}_{N}$ and by $X^{\mathrm{gen}}:=\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})^{\mathrm{gen}}$ for any $X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$.
@@
$X\mapsto X^{\mathrm{gen}}/\!\!/{N_{-}}:=(\mathfrak{n}_{-}\backslash X^{\mathrm{gen}})^{N_{-}}=\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}(\mathfrak{n}_{-}\backslash X)^{N_{-}},\ X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N}),$
@@
In fact, one can explicitly compute the parabolic restriction as follows. By [AST71], there exists an element $P$ of $\mathfrak{h}_{N}$-weight zero in a certain completion of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}[\hbar^{-1}]$ whose action is well-defined on any right $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}$-module $F$ with a locally-finite action of $\mathfrak{n}_{-}$ satisfying $F\mathfrak{n}_{-}\subset\hbar F$; this element is called the extremal projector. It satisfies the following properties:
(4.6) $P\in 1+\mathfrak{n}_{-}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\cap\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\mathfrak{n}_{+},\qquad E_{ij}P=PE_{ji}=0,\ 1\leq i<j\leq N,\qquad P^{2}=P.$
Notation 4.2.
@@
(4.7) $P\colon\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}/\mathfrak{n}_{+}\xrightarrow{\sim}(\mathfrak{n}_{-}\backslash X^{\mathrm{gen}})^{N_{-}},\qquad[x]\mapsto PxP.$
@@
(4.8) $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\otimes V\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes V/\!\!/N_{-},\qquad f(w)\otimes v\mapsto f(w)PvP,$
see [Zhe90].
@@
Usually, the cited results use the positive version of the parabolic restriction $X\mapsto(X^{\mathrm{gen}}/\mathfrak{n}_{+})^{N_{+}}$. For Miura transform of Section˜6, the negative one will be more convenient as the former is a map of left $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-modules; however, both versions are equivalent thanks to the property $\varpi_{N}(P)=P$. For instance, assume $\varpi_{N}$ can be extended to an isomorphism of a Harish-Chandra bimodule $X$. Then there is an anti-isomorphism
$\varpi_{N}\colon(\mathfrak{n}_{-}\backslash X^{\mathrm{gen}})^{N_{-}}\xrightarrow{\sim}(X^{\mathrm{gen}}/\mathfrak{n}_{+})^{N_{+}},\qquad PxP\mapsto\varpi_{N}(PxP)=P\varpi_{N}(x)P.$
@@
This is the standard Harish-Chandra isomorphism $\mathbf{Z}_{\mathfrak{gl}_{N}}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{W}$, see [Hum08, Theorem 1.10].
4.3. Vector representation
@@
(4.10) $\bar{v}_{i}=\sum\limits_{N\geq i_{1}>\ldots>i_{s}>i}P(w_{i}-w_{j_{1}})\ldots(w_{i}-w_{j_{r}})v_{i_{1}}E_{i_{2}i_{1}}\ldots E_{ii_{s}}\cdot\in\mathfrak{n}_{-}\backslash\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}.$
where $\{j_{1}\ldots j_{r}\}$ is the complementary subset to $\{i_{1},\ldots,i_{s}\}$ inside $\{i+1,\ldots,N\}$. According to formulas [Mol02, (3.2),(3.3)], the trick of Example˜3.6, and Remark˜4.3, those vectors are indeed invariant under $N_{-}$ and satisfy
(4.11) $\bar{v}_{i}=(w_{i}-w_{i+1})\ldots(w_{i}-w_{N})\cdot Pv_{i}P,$
@@
Using the parabolic restriction, one can deduce the following explicit description of the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$ as bimodule over the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$. Before the statement, we need the following lemma which is essentially a part of [KNV11, Theorem 2]; for reader’s convenience, we repeat its proof in our notations.
@@
$\mathrm{Hom}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes V,\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes W)\rightarrow\mathrm{Hom}_{\mathrm{HC}_{\hbar}(H_{N})^{\mathrm{gen}}}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes V/\!\!/N_{-},\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes W/\!\!/N_{-})$
@@
$s\in(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes W)^{\mathrm{GL}_{N}}/(\hbar)\cong(\mathcal{O}(\mathfrak{gl}_{N}^{*})\otimes W)^{\mathrm{GL}_{N}},$
@@
The equality in the following proposition is understood as in Remark˜4.8.
@@
For any $v\in\mathbf{C}^{N}$, we have
$A(u)^{-1}vA(u)=\frac{u-\Omega+\hbar}{u-\Omega}v.$
@@
Observe that both sides are actions of certain elements of $\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N})$. By ˜4.5, it is enough to show the equality under the parabolic restriction functor. It follows from ˜4.6 that
@@
At the same time, recall that by Example˜4.4, we have $PA(u)=\prod_{j=1}^{N}(u-w_{j})$, therefore,
$Pv_{i}A(u)P=P\prod_{j=1}^{N}v_{i}(u-w_{j})P=\prod_{j=1}^{N}(u-w_{i}+\delta_{ij}\hbar)\cdot Pv_{i}P$
@@
Strictly speaking, we understand the equality of Proposition˜4.7 in the sense of power series expansion in $u^{-1}$. The right-hand side can be presented as
$\frac{u-\Omega+\hbar}{u-\Omega}=1+\frac{\hbar}{u-\Omega}=1+\hbar\sum_{i=0}^{\infty}\Omega^{i}u^{-i-1}\in\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N})[[u^{-1}]],$
while the left-hand side can be equivalently presented as conjugation by $\mathrm{qdet}(T(-u+N\hbar-\hbar))$ in the notations of Section˜1.2 which also defines an element of $\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N})[[u^{-1}]]$. Another way is via the coefficients (1.11): for any $v\in\mathbf{C}^{N}$, we have
@@
In fact, using ˜4.6 and (4.9), one can re-deduce the Cayley-Hamilton theorem of Proposition˜3.7: observe that for every $i$, we have
@@
Then we apply ˜4.5.
4.4. Mickelsson algebras and difference operators
By [KS22, Corollary 4.18], the natural lax monoidal structure on $\mathrm{res}$ given by tensor multiplication
(4.12) $X^{\mathrm{gen}}/\!\!/N_{-}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}}Y^{\mathrm{gen}}/\!\!/N_{-}\rightarrow(X\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}Y)^{\mathrm{gen}}/\!\!/N_{-},\qquad[x]\otimes[y]\mapsto[x\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}}y]$
@@ 4.4. Mickelsson algebras and difference operators
Definition 4.10.
[Zhe90] Let $A\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$ be an algebra object. The Mickelsson algebra associated to $A$ is the $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodule $A^{\mathrm{gen}}/\!\!/N_{-}$ with the multiplication given by
$A^{\mathrm{gen}}/\!\!/N_{-}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}}A^{\mathrm{gen}}/\!\!/N_{-}\xrightarrow{\sim}(A^{\mathrm{gen}}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}}A^{\mathrm{gen}})/\!\!/N_{-}\rightarrow A^{\mathrm{gen}}/\!\!/N_{-},$
@@
$PaP\otimes PbP\mapsto PaPbP,\ a,b\in A^{\mathrm{gen}}.$
We will compute it explicitly for $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$. Consider the algebra $^{N}\tilde{\mathcal{A}}_{\hbar}=\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}[\mathbf{u}_{1}^{\pm},\ldots,\mathbf{u}_{N}^{\pm}]$ of difference operators with the relations
$\mathbf{u}^{\pm}_{i}f(w)=f(w\pm\hbar\epsilon_{i})\mathbf{u}_{i}^{\pm}.$
Denote by $^{N}\tilde{\mathcal{A}}_{\hbar}^{\pm}\subset{}^{N}\tilde{\mathcal{A}}_{\hbar}$ the $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-subalgebras generated by respectively positive or negative powers of $\{\mathbf{u}_{i}\}$.
@@
By (4.8), the monomials $Pv_{1}^{k_{1}}\ldots v_{N}^{k_{N}}P$ for all possible powers define a basis of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/N_{-}$ over $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$. By invertibility of the monoidal isomorphisms (4.12), the monomials
$(Pv_{1}P)^{k_{1}}\ldots(Pv_{N}P)^{k_{N}}$
for all possible powers also constitute a basis. Therefore, by (4.11), we conclude that, as a left $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$, the reduction $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/N_{-}$ is a free symmetric algebra in $\{\bar{v}_{i}\}$.
@@
where $E_{I}=E_{i_{2}i_{1}}\ldots E_{ii_{s}}\in\mathrm{U}_{\hbar}(\mathfrak{n}_{+})$ for $I=\{i_{1}>\ldots>i_{s}\}$. Consider the product $Pv_{i}Pv_{j}P$. By pushing terms of $\mathfrak{n}_{+}$ to the right using
@@
for all $a<b$, we conclude that the only terms contributing non-trivially are those with $I=\emptyset$ and $|I|=1$, so that
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/N_{-}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}[\bar{v}_{1},\ldots,\bar{v}_{N}]$
with relation (4.13). Also, it is clear that $^{N}\tilde{\mathcal{A}}_{\hbar}^{-}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}{}}[\mathbf{u}_{1}^{-1},\ldots,\mathbf{u}_{N}^{-1}]$ as left $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-modules. Therefore, the assignment $\bar{v}_{i}\mapsto\mathbf{u}^{-1}_{i}$ extends to an algebra isomorphism $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/N_{-}\xrightarrow{\sim}{}^{N}\tilde{\mathcal{A}}_{\hbar}^{-}$. ∎
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes S^{k}\mathbf{C}^{N}/\!\!/N_{-}\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}[\mathbf{u}_{1}^{-1},\ldots,\mathbf{u}_{N}^{-1}]_{k},$
where the target is the space of polynomials of total degree $k$ in $\mathbf{u}^{-1}$ variables. By ˜3.9, the source has an action of the spherical subalgebra $e^{\hbar}_{k}\mathfrak{H}^{\hbar}_{k}e^{\hbar}_{k}$ of the degenerate affine Hecke algebra.
@@
The spherical subalgebra $\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]^{S_{k}}$ acts on $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}[\mathbf{u}_{1}^{-1},\ldots,\mathbf{u}_{N}^{-1}]_{k}$ by
$f(\vec{Y})\cdot\mathbf{u}_{i_{1}}\ldots\mathbf{u}_{i_{k}}=\sum_{J}c_{J}(w_{i_{1}}^{j_{1}}\mathbf{u}^{-1}_{i_{1}})\cdot\ldots\cdot(w_{i_{k}}^{j_{k}}\mathbf{u}^{-1}_{i_{k}}),\qquad f(\vec{Y})=\sum_{J=(j_{1},\ldots,j_{k})}c_{J}Y_{1}^{j_{1}}\ldots Y_{k}^{j_{k}}$

### v4

@@ 5.2 Parabolic restriction
In particular, the tensor product $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}:=\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}\mathrm{U}_{\hbar}(\mathfrak{g})$ has a well-defined algebra structure. Denote by $\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})^{\mathrm{gen}}$ the corresponding generic version of Harish-Chandra bimodules for $\mathrm{GL}_{N}$ and by
$X^{\mathrm{gen}}:=\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})^{\mathrm{gen}}$
for any $X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$.
@@
$X\mapsto X^{\mathrm{gen}}/\!\!/{N_{-}}:=(\mathfrak{n}_{-}\backslash X^{\mathrm{gen}})^{N_{-}}=\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}(\mathfrak{n}_{-}\backslash X)^{N_{-}},\qquad X\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N}),$
@@
In fact, one can explicitly compute the parabolic restriction as follows. By [1], there exists an element $P$ of $\mathfrak{h}_{N}$-weight zero in a certain completion of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\bigl[\hbar^{-1}\bigr]$ whose action is well-defined on any right $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}$-module $F$ with a locally-finite action of $\mathfrak{n}_{-}$ satisfying $F\mathfrak{n}_{-}\subset\hbar F$; this element is called the extremal projector. It satisfies the following properties:
$P\in 1+\mathfrak{n}_{-}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\cap\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\mathfrak{n}_{+},$
$E_{ij}P=PE_{ji}=0,\qquad 1\leq i<j\leq N,$
$P^{2}=P.$ (5.5)
Notation 5.2.
@@
$P\colon\mathfrak{n}_{-}\backslash X^{\mathrm{gen}}/\mathfrak{n}_{+}\xrightarrow{\sim}(\mathfrak{n}_{-}\backslash X^{\mathrm{gen}})^{N_{-}},\qquad[x]\mapsto PxP.$ (5.6)
@@
$\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\otimes V\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes V/\!\!/N_{-},\qquad f(w)\otimes v\mapsto f(w)PvP,$ (5.7)
see [49].
@@
Usually, the cited results use the positive version of the parabolic restriction $X\mapsto(X^{\mathrm{gen}}/\mathfrak{n}_{+})^{N_{+}}$. For Miura transform of Section 7, the negative one will be more convenient as the former is a map of left $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-modules; however, both versions are equivalent thanks to the property $\varpi_{N}(P)=P$. For instance, assume $\varpi_{N}$ can be extended to an isomorphism of a Harish-Chandra bimodule $X$. Then there is an anti-isomorphism
[…]
@@
This is the standard Harish-Chandra isomorphism $\mathbf{Z}_{\mathfrak{gl}_{N}}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{W}$, see [25, Theorem 1.10].
5.3 Vector representation
@@
$\bar{v}_{i}=\sum\limits_{N\geq i_{1}>\dots>i_{s}>i}P(w_{i}-w_{j_{1}})\cdots(w_{i}-w_{j_{r}})v_{i_{1}}E_{i_{2}i_{1}}\cdots E_{ii_{s}}\cdot\in\mathfrak{n}_{-}\backslash\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N},$ (5.9)
where $\{j_{1},\dots,j_{r}\}$ is the complementary subset to $\{i_{1},\dots,i_{s}\}$ inside $\{i+1,\dots,N\}$. According to [37, formulas (3.2) and (3.3)], the trick of Example 4.6, and Remark 5.3, those vectors are indeed invariant under $N_{-}$ and satisfy
$\bar{v}_{i}=(w_{i}-w_{i+1})\cdots(w_{i}-w_{N})\cdot Pv_{i}P$ (5.10)
@@
Using the parabolic restriction, one can deduce the following explicit description of the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$ as bimodule over the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$. Before the statement, we need the following lemma which is essentially a part of [30, Theorem 2]; for reader’s convenience, we repeat its proof in our notations.
@@
$\mathrm{Hom}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes V,\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes W)$
$\qquad\rightarrow\mathrm{Hom}_{\mathrm{HC}_{\hbar}(H_{N})^{\mathrm{gen}}}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes V/\!\!/N_{-},\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes W/\!\!/N_{-})$
@@
$s\in(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes W)^{\mathrm{GL}_{N}}/(\hbar)\cong(\mathcal{O}(\mathfrak{gl}_{N}^{*})\otimes W)^{\mathrm{GL}_{N}}.$
@@
The equality in the following proposition is understood as in Remark 5.8.
@@
For any $v\in\mathbf{C}^{N}\subset\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$, we have
$A(u)^{-1}vA(u)=\frac{u-\Omega+\hbar}{u-\Omega}v$
as elements of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\bigl[\hskip-0.56905pt\bigl[u^{-1}\bigr]\hskip-0.56905pt\bigr]$.
@@
Both sides are actions of certain elements of $\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\bigr)\bigl[\hskip-0.56905pt\bigl[u^{-1}\bigr]\hskip-0.56905pt\bigr]$. By Lemma 5.5, it is enough to show the equality under their parabolic restriction functor. It follows from Lemma 5.6 that
@@
At the same time, recall that by Example 5.4, we have $PA(u)=\prod_{j=1}^{N}(u-w_{j})$, therefore,
$Pv_{i}A(u)P=P\prod_{j=1}^{N}v_{i}(u-w_{j})P=\prod_{j=1}^{N}(u-w_{i}+\delta_{ij}\hbar)\cdot Pv_{i}P.$
@@
Strictly speaking, we understand the equality of Proposition 5.7 in the sense of power series expansion in $u^{-1}$. The right-hand side can be presented as
$\frac{u-\Omega+\hbar}{u-\Omega}=1+\frac{\hbar}{u-\Omega}=1+\hbar\sum_{i=0}^{\infty}\Omega^{i}u^{-i-1}\in\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\bigr)\bigl[\hskip-0.56905pt\bigl[u^{-1}\bigr]\hskip-0.56905pt\bigr],$
while the left-hand side can be equivalently presented as conjugation by $\mathrm{qdet}(T(-u+N\hbar-\hbar))$ in the notations of Section 2.2 which also defines an element of $\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}\!\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\bigr)\!\bigl[\hskip-0.56905pt\bigl[u^{-1}\bigr]\hskip-0.56905pt\bigr]$. Another way is via the coefficients (2.8): for any $v\in\mathbf{C}^{N}$, we have
@@
In fact, using Lemma 5.6 and (5.8), one can re-deduce the Cayley–Hamilton theorem of Proposition 4.7: observe that for every $i$, we have
@@
Then we apply Lemma 5.5.
5.4 Mickelsson algebras and difference operators
By [28, Corollary 4.18], the natural lax monoidal structure on $\mathrm{res}$ given by tensor multiplication
$\begin{split}X^{\mathrm{gen}}/\!\!/N_{-}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}}Y^{\mathrm{gen}}/\!\!/N_{-}&\rightarrow(X\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}Y)^{\mathrm{gen}}/\!\!/N_{-},\\ [x]\otimes[y]&\mapsto\bigl[x\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}}y\bigr]\end{split}$ (5.11)
@@ 5.4 Mickelsson algebras and difference operators
Definition 5.10 ([49]).
Let $A\in\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})$ be an algebra object. The Mickelsson algebra associated to $A$ is the $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodule $A^{\mathrm{gen}}/\!\!/N_{-}$ with the multiplication given by
$A^{\mathrm{gen}}/\!\!/N_{-}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}}A^{\mathrm{gen}}/\!\!/N_{-}\xrightarrow{\sim}\bigl(A^{\mathrm{gen}}\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}}A^{\mathrm{gen}}\bigr)/\!\!/N_{-}\rightarrow A^{\mathrm{gen}}/\!\!/N_{-},$
@@
$PaP\otimes PbP\mapsto PaPbP,\qquad a,b\in A^{\mathrm{gen}}.$
We will compute it explicitly for $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$. Consider the algebra $^{N}\!\tilde{\mathcal{A}}_{\hbar}$ of (3.13). Denote by $^{N}\!\tilde{\mathcal{A}}_{\hbar}^{\pm}\subset{}^{N}\!\tilde{\mathcal{A}}_{\hbar}$ the $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-subalgebras generated by respectively positive or negative powers of $\{\mathbf{u}_{i}\}$.
@@
By (5.7), the monomials $Pv_{1}^{k_{1}}\cdots v_{N}^{k_{N}}P$ for all possible powers define a basis of the reduction $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/N_{-}$ over $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$. By invertibility of the monoidal isomorphisms (5.11), the monomials
$(Pv_{1}P)^{k_{1}}\cdots(Pv_{N}P)^{k_{N}}$
for all possible powers also constitute a basis. Therefore, by (5.10), we conclude that, as a left $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-module, the reduction $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/N_{-}$ is a free symmetric algebra in $\{\bar{v}_{i}\}$.
@@
where $E_{I}=E_{i_{2}i_{1}}\cdots E_{ii_{s}}\in\mathrm{U}_{\hbar}(\mathfrak{n}_{+})$ for $I=\{i_{1}>\dots>i_{s}\}$. Consider the product $Pv_{i}Pv_{j}P$. By pushing terms of $\mathfrak{n}_{+}$ to the right using
@@
for all $a<b$, we conclude that the only terms contributing non-trivially are those with $I=\varnothing$ and $|I|=1$, so that
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/N_{-}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}[\bar{v}_{1},\dots,\bar{v}_{N}]$
with relation (5.12). Also, it is clear that $^{N}\!\tilde{\mathcal{A}}_{\hbar}^{-}\cong\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}{}}\bigl[\mathbf{u}_{1}^{-1},\dots,\mathbf{u}_{N}^{-1}\bigr]$ as left $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-modules. Therefore, the assignment $\bar{v}_{i}\mapsto\mathbf{u}^{-1}_{i}$ extends to an algebra isomorphism
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\ltimes S^{\bullet}\mathbf{C}^{N}/\!\!/N_{-}\xrightarrow{\sim}{}^{N}\!\tilde{\mathcal{A}}_{\hbar}^{-},$
as claimed. ∎
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes S^{k}\mathbf{C}^{N}/\!\!/N_{-}\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\bigl[\mathbf{u}_{1}^{-1},\dots,\mathbf{u}_{N}^{-1}\bigr]_{k},$
where the target is the space of polynomials of total degree $k$ in $\mathbf{u}^{-1}$ variables. By Corollary 4.9, the source has an action of the spherical subalgebra ${\rm e}^{\hbar}_{k}\mathfrak{H}^{\hbar}_{k}{\rm e}^{\hbar}_{k}$ of the degenerate affine Hecke algebra.
@@
The spherical subalgebra $\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]^{S_{k}}$ acts on localized difference operators $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\bigl[\mathbf{u}_{1}^{-1},\dots,\mathbf{u}_{N}^{-1}\bigr]_{k}$ by
$f\bigl(\vec{Y}\bigr)\cdot\mathbf{u}_{i_{1}}^{-1}\cdots\mathbf{u}_{i_{k}}^{-1}=\sum_{J}c_{J}\bigl(w_{i_{1}}^{j_{1}}\mathbf{u}^{-1}_{i_{1}}\bigr)\cdots(w_{i_{k}}^{j_{k}}\mathbf{u}^{-1}_{i_{k}}),\qquad f\bigl(\vec{Y}\bigr)=\sum_{J=(j_{1},\dots,j_{k})}c_{J}Y_{1}^{j_{1}}\cdots Y_{k}^{j_{k}}$